EDITORIAL

Terry Goodman

With my summer schedule allowing for a little more “free” time, I was able to complete my reading of The Teaching Gap, by James W. Stigler and James Hiebert (The Free Press, 1999). Using a series of videos that were produced from the Third International Mathematics and Science Study (TIMSS), these researchers compare mathematics teaching practices in Japan and Germany with those in the United States. I found the comparisons to be quite striking and revealing.

Stigler and Hiebert suggest that the task of improving American students’ performance in mathematics will require solutions that focus on impacting the quality of teaching inside classrooms. Teaching, they argue, is cultural. They propose that, not teachers, but patterns of teaching must be changed. They further propose ways schools can be restructured so that teachers can engage in collaborative lesson study and development. I encourage you to read this book.

Once again, we would like to acknowledge the contributions our referees make to the quality of MJMS. We deeply appreciate their time and effort.

List of Referees

Hadi Y. Alkahby, Dillard University Charles J. Allard, Polo R-VII High School Mohammad K. Azarian, University of Evansville Seung–Jin Bang, Seoul, Korea Gerald Bergum, South Dakota State University John E. Bernard, University of Texas–Pan American Jim Bruening, Southeast Missouri State University Joseph Chance, University of Texas–Pan American Hang Chen, Central Missouri State University Eungchun Cho, Kentucky State University Gregory Davis, University of Wisconsin–Green Bay Bryan Dawson, Union University Richard Delaware, University of Missouri–Kansas City Joseph B. Dence, University of Missouri–St. Louis Albert Dixon, College of the Ozarks Clayton Dodge, University of Maine John J. Edgell, Jr., Southwest Texas State University Vince Edmondson, Central Missouri State University

List of Referees (cont.)

Larry Eifler, University of Missouri–Kansas City Ekemezie J. Emeka, Columbia College Martin Erickson, Truman State University Russell Euler, Northwest Missouri State University Joe Flowers, Trinity University Jayanthi Ganapathy, University of Wisconsin–Oshkosh Ramesh Garimella, Northwest Missouri State University Raymond J. Grinnell, University of the West Indies Victor Gummersheimer, Southeast Missouri State University K.L.D. Gunawardena, University of Wisconsin–Oshkosh Dennis R. Harmon, Missouri Southern State College Stan Hartzler, Virginia Polytech Institute & State University Troy Hicks, University of Missouri–Rolla Joe Howard, New Mexico Highlands University Timothy P. Keller, George Mason University N. J. Kuenzi, University of Wisconsin–Oshkosh Charles E. Lamb, University of Texas–Austin Mulatu Lemma, Savannah State University Man Chun Leung, National University of Singapore Yi–Hsin Liu, University of Nebraska–Omaha William Livingston, Missouri Southern State College Roger Marty, Cleveland State University Wayne McDaniel, University of Missouri–St. Louis Gary McDonald, Northwest Missouri State University Rhonda L. McKee, Central Missouri State University Phoebe McLaughlin, Central Missouri State University Peter R. Mercer, St. Mary’s College of Maryland Kandasamy Muthvel, University of Wisconsin–Oshkosh Ollie Nanyes, Bradley University Krzysztof Ostaszewski, University of Louisville Miguel Paredes, University of Texas–Pan American Don Redmond, Southern Illinois University Jeffrey A. Rosoff, Gustavus Adolphus College Prasanna Sahoo, University of Louisville Jau–Shyong Shiue, University of Nevada–Las Vegas Donald P. Skow, University of Texas–Pan American Randall J. Swift, Western Kentucky University Jinhua Tao, Central Missouri State University Songlin Tian, Central Missouri State University Mohan Tikoo, Southeast Missouri State University Mathew Timm, Bradley University Alvin Tinsley, Central Missouri State University Peter L. Vachuska, University of Wisconsin–Washington Co. Bobby Neal Winters, Pittsburg State University Fred Worth, Henderson State University Joseph Wiener, University of Texas–Pan American